OpenStudy (anonymous):

how to do the limit as x approaches 0 of (sin4x)/(tan5x)

OpenStudy (jamesj):

If you know l'Hopital's rule, this wouldn't be a bad place to use it. If not or if you want to avoid the careful writing out of l'Hopital rule application, then write \[\frac{\sin 4x}{\tan 5x} = \frac{4}{5} \frac{(\sin 4x)/ 4x}{(\tan 5x)/5x}\] Then the limit of that last fraction is 1/1